Circular No. 6573

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) D. C. Lis, M. Gardner, and T. G. Phillips, Caltech Submillimeter Ob-

servatory (CSO); D. Bockelée-Morvan, N. Biver, J. Crovisier, H. Rauer, P. Colom, and D. Gautier, Observatoire de Paris; and D. Despois, Observatoire de Bordeaux, report the radio detection of sulfur monoxide (SO), carbonyl sulfide (OCS), ionized carbon monoxide (CO^{+}), and the cyanogen radical (CN) at the CSO on Feb. 20–23: “The N(J)=5(5)–5(4) 215.22-GHz, 5(6)–4(5) 251.87-GHz, and 8(7)–7(6) 304.08-GHz transitions of SO were detected with integrated line areas on a T_{b} scale of 0.11 ± 0.02, 0.33 ± 0.04, and 0.50 ± 0.05 K km/s, respectively. This is the first secure detection of SO in a comet. OCS was detected through its J(25–24) line at 303.99 GHz (0.15 ± 0.03 K km/s ). We have also detected for the first time rotational transitions of CO^{+} [N(J)=2(5/2)–1(3/2) at 236.06 GHz] and CN [2(5/2)–1(3/2) at 226.87 GHz] with line areas of 0.43 ± 0.03 and 0.21 ± 0.04 km/s, respectively. The OCS production rate is estimated to be 1.3 × 10^{28} molecules/s. The 252-GHz lines of methanol show a rotational temperature of 70 ± 3 K.”

M. J. Mumma, M. A. DiSanti, and N. Dello Russo, Goddard Space Flight Center, NASA; K. Magee-Sauer, Rowan College of New Jersey; R. Novak, Iona College; and M. Fomenkova, University of California at San Diego, report: “We detected parent volatiles in C/1995 O1 on Feb. 23.9– 24.1 and Mar. 1.7–1.9 UT, using the NASA Infrared Telescope Facility (+ CSHELL cryogenic infrared spectrometer) on Mauna Kea. The species, the vibrational bands, and the detected lines are: CO [1-0, 10 lines, J′ = 0–9]; C_{2}H_{6} [ν_{7}, eight Q branches]; CH_{3}OH ν_{3} Q-branch; numerous lines of H_{2}O and of HDO [ν_{1}]. HCN, C_{2}H_{2}, and CH_{4} were also detected. All lines are peaked on the nucleus, and most are extended; CO extends to the ends of the east-west slit, ∼ 15′′ in either direction. For an aperture 1′′ × 1′′.4 centered on the nucleus, we obtained the following line fluxes: CO R3, 26 × 10^{−17} W m^{−2}; C_{2}H_{6} ^{r}Q_{0}-branch, 3.2 × 10^{−17} W m^{−2}, ^{r}Q_{1}- branch, 4.5 × 10^{−17} W m^{−2}; CH_{4} ν_{3} R0, 2.8 × 10^{−17} W m^{−2}. The flux density for the continuum is 1.8 × 10^{−15} W m^{−2} cm^{−1} at 3.35 μm. The CO R3 line is enhanced by a factor of three, compared with Jan. 21.5. The dust continuum is markedly stronger than on Jan. 21.5, and its spectrum suggests very strong emission from organic grains. A preliminary analysis of C_{2}H_{6} ν_{7} suggests a rotational temperature in the range 60–110 K.”

Naked-eye m_{1} and tail-length estimates by M. V. Zanotta, near Milan, Italy: Feb. 28.19 UT, 0.6, 6^{o}; Mar. 2.19, 0.4, 8^{o}; 3.19, 0.3, 9^{o}.

1997 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green